demand: Q=25-0.5P supply: Q=10+P find the equillibrium price and quantity


1
Expert's answer
2020-07-30T11:47:54-0400
"Q_D=25-0.5p"

"Q_S=10+p"


"Q_D=Q_S"

"25-0.5p=10+p"

"1.5p=15"

"p=10"

"Q=20"